OpenStudy (anonymous):

Actually, my question is about physical-chemistry... How do I calculate the middle activitycoefficient?

OpenStudy (anonymous):

\[\ln f \pm = \ln 10\times logf \pm=\frac{ F }{ 2RT }\left( E ^{0_{Ag|AgCl|Cl ^{-} }}- \Delta E _{Pt|H _{2}|HCl|AgCl|Ag}\right)- \ln c _{HCl}\]

OpenStudy (anonymous):

So my question would be, how to derrive the single E's? The concentration of HCl is 5x10^-4 m.